OpenStudy (anonymous):

Find all solutions in the interval [0, 2pi) (Give me time to write it out.)

OpenStudy (anonymous):

\[\sin x \div 2 + cosx -1 =0\]

OpenStudy (anonymous):

\[\sin(\frac{x}{2}) + \cos(x) -1=0\] ?

OpenStudy (anonymous):

Yeah, how do you make the fraction bar like that? I can only seem to make the 2nd grade one.

OpenStudy (anonymous):

In the equation editor, type frac{numerator}{denominator}

OpenStudy (anonymous):

Okay\[\frac{4}{3}\] Sweet.

OpenStudy (anonymous):

But back to the question!

OpenStudy (anonymous):

I'm gonna mull it over while eatin' dinner. If it's not answered when I come back, I'll give it a shot.

OpenStudy (anonymous):

Okay! I had cheeseburgers for dinner, just so you're informed...

OpenStudy (anonymous):

I'm jealy.

OpenStudy (anonymous):

I know. But go enjoy your dinner.